A 5.00 L of a gas is collected at 100. K and then allowed to expand to 20.0 L. What must the new temperature be in order to maintain the same pressure (as required by Charles' Law)?

1
Expert's answer
2022-03-30T02:37:01-0400

Charles' Law:

V1/T1 = V2/T2

We have 3 of the 4 variables, and if we fill those in, we have...

5.00/100 = 20.0/T2

Using the algebra skill of cross multiplying, we can get...

5.00T2 = 2000

Dividing both sides by 5.00 to isolate the T2, we get...

T2 = 400k